OpenStudy (anonymous):

solve and write the solution in interval notation: |2b-7| is greater than or equal to 9

OpenStudy (zzr0ck3r):

we have |2b-7| >= 9 split it into two equations 2b-7>=9 2b-7<= -9 solve for b in both equations